OpenStudy (anonymous):

A particle moves on a straight line v(t)=sinwt cos ^{3} wt. find the position s(t) if s(0)=0. I've done this style of question before, but the two variables throw me off. I assume that u=cos(wt) but am unsure.

OpenStudy (anonymous):

then du=-sin(wt)

OpenStudy (anonymous):

-du=sin(wt)

OpenStudy (anonymous):

\[- \int\limits_{?}^{?} (u)^{3} du\]

OpenStudy (anonymous):

\[-\frac{ 1 }{ 4 } (u)^{4} +c\]

OpenStudy (anonymous):

\[- \frac{ (coswt)^{4} }{ 4 } +c\] then insert s(0)=0 so..... \[- \frac{ (cosw(0))^{4} }{ 4 } +c\]=0 ... but w is still an unknown...

OpenStudy (anonymous):

woooooooo no idea whats going on....

OpenStudy (anonymous):

unless -cosw(0)=0... then c=0... yeah i'm confused, doesn't seem right.

OpenStudy (sirm3d):

\[\mathrm du=-w\cdot\sin wt\,\mathrm dt\]